# Adaptive Uncertainty Quantification for Trajectory Prediction Under Distributional Shift

### v1

@@ Abstract
Trajectory prediction models that can infer both finite future trajectories and their associated uncertainties of the target vehicles in an online setting (e.g., real-world application scenarios) is crucial for ensuring the safe and robust navigation and path planning of autonomous vehicle motion. However, the majority of existing trajectory prediction models have neither considered reducing the uncertainty as one objective during the training stage nor provided reliable uncertainty quantification during inference stage under potential distribution shift. Therefore, in this paper, we propose the Conformal Uncertainty Quantification under Distribution Shift framework, CUQDS, to quantify the uncertainty of the predicted trajectories of existing trajectory prediction models under potential data distribution shift, while considering improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. Specifically, CUQDS includes 1) a learning-based Gaussian process regression module that models the output distribution of the base model (any existing trajectory prediction or time series forecasting neural networks) and reduces the estimated uncertainty by additional loss term, and 2) a statistical-based Conformal P control module to calibrate the estimated uncertainty from the Gaussian process regression module in an online setting under potential distribution shift between training and testing data. Experimental results on various state-of-the-art methods using benchmark motion forecasting datasets demonstrate the effectiveness of our proposed design.
@@ 1 Introduction
Accurately and efficiently predicting the future trajectories of the target vehicles is pivotal for ensuring the safety of path planning and navigation of autonomous systems in dynamic environments Hu et al. (2023); Kedia et al. (2023). However, trajectory uncertainly due to the changing external surrounding environments (e.g., road networks, neighborhood vehicles, passengers, and so on) or intrinsic intention changes of drivers, can lead to both trajectory distribution shift that change over time and overconfidence in the output of trajectory prediction models. It remains challenging to predict the distribution of the future trajectories of the target vehicles (rather than focusing solely on point estimates), and to quantify the uncertainty of the output trajectories under potential distribution shift between training ans testing data.
@@
    We propose CUQDS framework that integrates both the learning-based and statistical-based modules to provide uncertainty quantification for time series trajectory prediction under distribution shift. The proposed learning-based module is trained alongside the base model to quantify its output uncertainty. In the training stage, the main objectives are to enhance prediction accuracy of base model and reduce the estimated uncertainty by incorporating an additional loss term.
@@ 3.1 Problem Formulation
We assume that we have a time series trajectory prediction model $f$ with parameters $\theta$. We call this model $f_{\theta}$ a base model and it can be implemented as different structures of neural network Zhou et al. (2022); Liu et al. (2021); Zhou et al. (2023); Liang et al. (2020); Zhong et al. (2022). The model $f_{\theta}$ is trained using the dataset $\mathcal{D}$, where the target trajectory, denoted as $Y$, is conditioned on input $x$ and the dataset $\mathcal{D}$. We model the conditional distribution of $Y$ given $x$ and $\mathcal{D}$ as a Gaussian distribution, where $Y|x,\mathcal{D}\sim\mathcal{N}\left(\mu(x),\sigma^{2}(x)\right)$. In this formulation, $\mu(x)$ and $\sigma^{2}(x)$ are functions that map the input $x$ to the mean and variance of the Gaussian distribution, respectively. We treat the predicted trajectory of the base model $f_{\theta}$ as the mean $\mu(x)$ directly. Our main task in this study is to estimate the variance $\sigma^{2}(x)$ to denote the output uncertainty.
We propose the Conformal Uncertainty Quantification framework, CUQDS, as illustrated in Fig. 2, to quantify the uncertainty of the predicted trajectory of base model under potential distribution shift. The main objectives of this framework are to improve prediction accuracy of base model and reduce the output uncertainty. The framework introduces two modules, 1) In particular, in the training and validation stages ( Alg. 1), we propose a learning-based Gaussian process regression module to approximate the conditional output distribution $(Y|x,\mathcal{D})$ of base model by estimating $(Y|x,\theta\sim\mathcal{N}\left(\hat{\mu}\left(x\right),\hat{{\sigma}}^{2}\left(x\right)\right))$ based on $f_{\theta}$. The details are introduced in Sec. 3.2. 2) In the testing or inference stage (Algorithm. 2), we propose a statistical-based conformal P control module to calibrate the output distribution $\tilde{Y}|\tilde{x},\theta\sim\mathcal{N}\left(\tilde{\mu}\left(\tilde{x}\right),\tilde{{\sigma}}^{2}\left(\tilde{x}\right)\right)$ by considering the potential distribution shift between the training and testing datasets and the performance limitation of $f_{\theta}$ on current time series trajectory input. The conformal quantile of the module is first initialized using the validation data and keep updating after each prediction during the inference stage. We build the calibrated uncertainty interval and ensure it covers the true target trajectory with a predefined probability in long-run. The details are introduced in Sec. 3.3.
@@
where $w_{1}\in\mathbb{R}$ and $w_{2}\in\mathbb{R}$ are the weights adjusting the influence of two loss terms, respectively. During testing or inference stage, we will calibrate the covariance and estimate the uncertainty interval of the predicted trajectory.
@@ 3.2 Gaussian Process Regression Module
Existing literature of trajectory prediction models Zhou et al. (2022, 2023); Liang et al. (2020); Zhong et al. (2022) focuses on providing point estimates of future trajectories. However, the uncertainty of the future trajectories due to the changing environment or the intrinsic intention changes of drivers can lead to significant distribution shift and overconfident trajectory prediction. Such distribution shift and overconfident prediction can greatly impact the subsequent decision-making processes, such as robust path planning Hu et al. (2023); Kedia et al. (2023). Hence, we propose to consider output uncertainties from both the prediction capability limitation of the model $f_{\theta}$ upon current trajectory inputs and the noise inherent in the trajectory data.
@@ 3.2.4 Inference under Distribution Shift
Through the Gaussian process regression with a learnable kernel function, we are able to estimate the target trajectory’s distribution that in the training/validation dataset. However, under potential distribution shift between training and testing data in real-world settings, solely keep using the information learn from the training/validation datasets will lead to high generalization error. To take into account the potential distribution shift between the dataset $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2}\}$ and the testing dataset $\mathcal{\tilde{D}}$, we instead estimate the conditional output distribution of base model in the testing stage by $\tilde{Y}|\tilde{x},\mathcal{D}\sim\mathcal{N}(\tilde{\mu}(\tilde{x}),\tilde{\sigma}(\tilde{x}))$, where $(\tilde{x},\tilde{y})\in\tilde{\mathcal{D}}$. Then we compute the variance of the conditional output distribution of base model in the testing stage by
@@
To find the optimal parameters for the covariance function and the noise $\epsilon$, we introduce a new loss term $\mathcal{L}_{2}$:
@@ 3.3 Uncertainty Quantification under Distribution Shift through Calibration
The learning-based Gaussian process regression module is prone to overfit the training data and provide overconfident uncertainty estimation under distribution shift. Moreover, such uncertainty is more concerned about the aleotoric uncertainty, referring to the probabilistic nature of noise in data Kendall and Gal (2017), but fails to consider the performance limitation of base model on different trajectory data, namely the epistemic uncertainty. To solve these problems, we propose a conformal P control module to calibrate the output uncertainty from the Gaussian process regression module by considering the performance of base model under potential distribution shift. The idea of this method is based on both the conformal prediction Shafer and Vovk (2008) and the P control in Angelopoulos et al. (2023). However, different from standard conformal prediction, we assume the training data and testing data follow different distributions. We initialize the conformal quantile using the validation data and keep updating it after every prediction step in the inference stage to provide trustworthy uncertainty quantification.
3.3.1 Conformal Prediction for Long-run Converage
@@ 4.1 Experimental Setups
Uncertainty Evaluation Metrics: To further verify the efficacy of our CUQDS in reducing the predicted uncertainty to provide narrow while accurate uncertainty interval, we adopt the Negative Log-Likelihood (NLL) Feng et al. (2021) to assess the level of uncertainty in the predicted distribution. For this metric, lower values connote a higher degree of precision in uncertainty estimation and narrower uncertainty interval.
@@ 4.2 Baselines
Table 3: Compare the coverage rates of the uncertainty interval with and without the uncertainty calibration (UC) module.
@@ 4.4 Ablation Study
$\bullet$ Conformal Prediction VS conformal P control module: We conduct ablation study on replacing the P control uncertainty calibration module with the standard split conformal prediction Shafer and Vovk (2008). More specificity, the standard split conformal prediction (CP) estimates the conformal quantile $\hat{q}$ as the $\frac{(1-\alpha)(N_{2}+1)}{N_{2}}$ smallest element in the conformal score set $\mathcal{S}={\{s\left(x_{i},y_{i}\right)\}}_{i=1}^{N_{2}}$ which uses validation dataset. This conformal quantile $\hat{q}$ is only calculated once and fixed in the testing stage to calibrate the uncertainty.
@@
$\bullet$ Gaussian process regression module VS direct modeling: To validate the effectiveness of our Gaussian process regression module, we replace the Gaussian process regression module with the self-attention based design Mao et al. (2023) to predict the variance of the output distribution and add the KLD loss term to reduce the estimated uncertainty during the training stage. We apply this setting on the base model HiVT. As shown in Table. 4, both of the methods achieves better prediction accuracy than without considering uncertainty in the base model. Our CUQDS slightly exceed the direct modeling method in all prediction accuracy evaluation metrics. The results prove that our CUQDS is sufficient in providing output uncertainty estimation for the trajectory prediction base model and improving the prediction accuracy.
@@ 5 Conclusion & Discussion
In this study, we present the framework CUQDS to estimate the output distribution for any existing trajectory prediction models under distribution shift, while improving the prediction accuracy and reduce the output uncertainty by introducing additional loss term. CUQDS adopts the Gaussian process regression module to model the output distribution of the base model. In addition, CUQDS utilizes a statistical-based P control module to calibrate the estimated uncertainty by considering the performance limitation of base model upon recent inputs. The experiment results demonstrate the efficacy of CUQDS in improving the prediction accuracy and reducing the prediction uncertainty. Our findings highlight the importance of quantifying the uncertainty of output trajectory in trajectory prediction under distribution shift. In the future work, we plan to extend our work to more state-of-the-art models, provide uncertainty of output trajectory in each time step, and further compare the performance difference of estimating uncertainty and output distribution by different designs.

### v2

# CUQDS: Conformal Uncertainty Quantification under Distribution Shift for Trajectory Prediction

@@ Abstract
Trajectory prediction models that can infer both future trajectories and their associated uncertainties of the target vehicles is crucial for safe and robust navigation and path planning of autonomous vehicles. However, the majority of existing trajectory prediction models have neither considered reducing the uncertainty as one objective during the training stage nor provided reliable uncertainty quantification during inference stage, especially under potential distribution shift. Therefore, in this paper, we propose the Conformal Uncertainty Quantification under Distribution Shift framework, CUQDS, to quantify the uncertainty of the predicted trajectories of existing trajectory prediction models under potential data distribution shift, while improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. Specifically, CUQDS includes 1) a learning-based Gaussian process regression module that models the output distribution of the base model (any existing trajectory prediction neural networks) and reduces the estimated uncertainty by an additional loss term, and 2) a statistical-based Conformal P control module to calibrate the estimated uncertainty from the Gaussian process regression module in an online setting under potential distribution shift between training and testing data. Experimental results on various state-of-the-art methods using benchmark motion forecasting datasets demonstrate the effectiveness of our proposed design.
@@ 1 Introduction
Accurately and efficiently predicting the future trajectories of the target vehicles is pivotal for ensuring the safety of path planning and navigation of autonomous systems in dynamic environments Hu et al. (2023); Kedia et al. (2023). However, trajectory uncertainly due to the changing external surrounding environments (e.g., road networks, neighborhood vehicles, passengers, and so on) or intrinsic intention changes of drivers, can lead to both trajectory distribution shift that change over time and overconfidence in the output of trajectory prediction models. It remains challenging to predict the distribution of the future trajectories of the target vehicles (rather than focusing solely on point estimates), and to quantify the uncertainty of the output trajectories under potential distribution shift between training and testing data.
@@
    We propose CUQDS framework that integrates both the learning-based and statistical-based modules to provide uncertainty quantification for trajectory prediction under distribution shift. The proposed learning-based module is trained alongside the base model to quantify its output uncertainty. In the training stage, the main objectives are to enhance prediction accuracy of base model and reduce the estimated uncertainty by incorporating an additional loss term.
@@ 3.1 Problem Formulation
We assume that we have a trajectory prediction model $f$ with parameters $\theta$. We call this model $f_{\theta}$ a base model and it can be implemented as different structures of deep neural network Zhou et al. (2022); Liu et al. (2021); Zhou et al. (2023); Liang et al. (2020); Zhong et al. (2022). The model $f_{\theta}$ is trained using the dataset $\mathcal{D}$, where the target trajectory, denoted as $Y$, is conditioned on input $x$ and the dataset $\mathcal{D}$. We model the conditional distribution of $Y$ given $x$ and $\mathcal{D}$ as a Gaussian distribution, where $Y|x,\mathcal{D}\sim\mathcal{N}\left(\mu(x),\sigma^{2}(x)\right)$. In this formulation, $\mu(x)$ and $\sigma^{2}(x)$ are functions that map the input $x$ to the mean and variance of the Gaussian distribution, respectively. We treat the predicted trajectory of the base model $f_{\theta}$ as the mean $\mu(x)$ directly. Our main task in this study is to estimate the variance $\sigma^{2}(x)$ to denote the output uncertainty.
We propose the Conformal Uncertainty Quantification framework, CUQDS, as illustrated in Fig. 2, to quantify the uncertainty of the predicted trajectory of base model under potential distribution shift. The main objectives of this framework are to improve prediction accuracy of base model and reduce the output uncertainty. The framework introduces two modules, 1) In particular, in the training and validation stages ( Alg. 1), we propose a learning-based Gaussian process regression module to approximate the conditional output distribution $(Y|x,\mathcal{D})$ of base model by estimating $\left(Y|x,\theta\sim\mathcal{N}\left(\hat{\mu}\left(x\right),\hat{{\sigma}}^{2}\left(x\right)\right)\right)$ based on $f_{\theta}$. 2) In the testing or inference stage (Algorithm. 2), we propose a statistical-based conformal P control module to calibrate the output distribution $\tilde{Y}|\tilde{x},\theta\sim\mathcal{N}\left(\tilde{\mu}\left(\tilde{x}\right),\tilde{{\sigma}}^{2}\left(\tilde{x}\right)\right)$ by considering the potential distribution shift between the training and testing datasets and the performance limitation of $f_{\theta}$ on current time series trajectory input. The conformal quantile of the module is first initialized using the validation data and keep updating after each prediction during the inference stage. We build the calibrated uncertainty interval and ensure it covers the true target trajectory with a predefined probability in long-run.
@@
where $w_{1}\in\mathbb{R}$ and $w_{2}\in\mathbb{R}$ are the weights adjusting the influence of two loss terms, respectively. During inference stage, we will calibrate the estimated uncertainty of the predicted trajectory.
@@ 3.2 Gaussian Process Regression Module
Existing literature of trajectory prediction models Zhou et al. (2022, 2023); Liang et al. (2020); Zhong et al. (2022) focuses on providing point estimates of future trajectories. However, the uncertainty of the future trajectories due to the changing environment or the intrinsic intention changes of drivers can lead to significant distribution shift and overconfident trajectory prediction. Such distribution shift and overconfident prediction can greatly impact the subsequent decision-making processes, such as robust path planning Hu et al. (2023); Kedia et al. (2023). Hence, we propose to consider output uncertainties from both the modeling limitation of the model $f_{\theta}$ upon current trajectory inputs and the noise inherent in the trajectory data.
@@ 3.2.4 Inference under Distribution Shift
Through the Gaussian process regression with a learnable kernel function, we are able to estimate the target trajectory’s distribution that in the training/validation dataset. However, under potential distribution shift between training and testing data in real-world settings, solely keep using the information learn from the training/validation datasets will lead to high generalization error. To take into account the potential distribution shift between the dataset $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2}\}$ and the testing dataset $\mathcal{\tilde{D}}$, we instead estimating the conditional output distribution of base model in the testing stage by $\tilde{Y}|\tilde{x},\mathcal{D}\sim\mathcal{N}(\tilde{\mu}(\tilde{x}),\tilde{\sigma}(\tilde{x}))$, where $(\tilde{x},\tilde{y})\in\tilde{\mathcal{D}}$. Then we compute the variance of the conditional output distribution of base model in the testing stage by
@@
To find the optimal parameters for the covariance function and the noise $\epsilon$, and reduce the estimated uncertainty, we introduce a new loss term $\mathcal{L}_{2}$:
@@ 3.3 Uncertainty Quantification under Distribution Shift through Calibration
The learning-based Gaussian process regression module is prone to overfit the training data and provide overconfident uncertainty estimation under distribution shift. Moreover, such uncertainty is more concerned about the aleatoric uncertainty, referring to the probabilistic nature of noise in data Kendall and Gal (2017), but fails to consider the performance limitation of base model on different trajectory data, namely the epistemic uncertainty. To solve these problems, we propose a conformal P control module to calibrate the output uncertainty from the Gaussian process regression module by considering the performance of base model under potential distribution shift. The idea of this method is based on both the conformal prediction Shafer and Vovk (2008) and the P control in Angelopoulos et al. (2023). However, different from standard conformal prediction, we assume the training data and testing data follow different distributions. We initialize the conformal quantile using the validation data and keep updating it after every prediction step in the inference stage to provide trustworthy uncertainty quantification.
3.3.1 Conformal Prediction for Long-run Coverage
@@ 4.1 Experimental Setups
Uncertainty Evaluation Metrics: To further verify the efficacy of our CUQDS in reducing the predicted uncertainty to provide narrow while accurate uncertainty interval, we adopt the Negative Log-Likelihood (NLL) Feng et al. (2021) to assess the level of uncertainty in the predicted distribution. Lower values connote a higher degree of precision in uncertainty estimation and narrower uncertainty interval.
@@ 4.2 Baselines
Table 3: The coverage rates of the uncertainty interval with and without the uncertainty calibration (UC) module.
@@ 4.4 Ablation Study
$\bullet$ Conformal Prediction VS conformal P control module: We conduct ablation study on replacing the P control uncertainty calibration module with the standard split conformal prediction Shafer and Vovk (2008). More specificity, the standard split conformal prediction (CP) estimates the conformal quantile $\hat{q}$ as the $\frac{(1-\alpha)(N_{2}+1)}{N_{2}}$ smallest element in the conformal score set $\mathcal{S}={\{s\left(x_{i},y_{i}\right)\}}_{i=1}^{N_{2}}$ from validation dataset. This conformal quantile $\hat{q}$ is only calculated once and fixed in the testing stage to calibrate the uncertainty.
@@
$\bullet$ Gaussian process regression module VS direct modeling: To validate the effectiveness of our Gaussian process regression module, we replace the Gaussian process regression module with the self-attention based design Mao et al. (2023) to predict the variance of the output distribution and add the KLD loss term to reduce the estimated uncertainty during the training stage. We apply this setting on the base model HiVT. As shown in Table. 4, both of the methods achieve better prediction accuracy than without considering uncertainty in the base model. Our CUQDS slightly exceed the direct modeling method in all prediction accuracy evaluation metrics. The results prove that our CUQDS is sufficient in providing output uncertainty estimation for the trajectory prediction base model and improving the prediction accuracy.
@@ 5 Conclusion & Discussion
In this study, we present the framework CUQDS to estimate the output distribution for any existing trajectory prediction models under distribution shift, while improving the prediction accuracy and reducing the output uncertainty. CUQDS adopts the Gaussian process regression module to model the output distribution of the base model. In addition, CUQDS utilizes a statistical-based P control module to calibrate the estimated uncertainty by considering the performance limitation of base model upon recent inputs. The experiment results demonstrate the efficacy of CUQDS in improving the prediction accuracy and reducing the prediction uncertainty. Our findings highlight the importance of quantifying the uncertainty of output trajectory in trajectory prediction under distribution shift. In the future work, we plan to extend our work to more state-of-the-art models, provide uncertainty of output trajectory in each time step, and further compare the performance difference of estimating uncertainty and output distribution by different designs.